Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2000-253_2000-259

Mission Operations Days: 2000/253 to 2000/259
September 09, 2000 00:00:00 GMT to September 15, 2000 00:00:00 GMT


NOTE: Band 14H (detectors 1-10) and Band 25 (detector 1) gain changes were made on day 256 (Sept. 12th) at approx. 18:17:46 GMT.


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
ComponentState(s)Comment
BlackbodyA On (290 K); B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On, #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA On; B OffNominal
Flight SoftwareRev BD + 3 sets of patches
  • FR Error Counting and Statistics (2 patches)
  • FR Band 14H Gain Change (10 commands)
  • FR Band 25 Gain Change (Detector #1) (1 command)
Nominal
Inhibit Ids SetNoneNominal
TMONs Enabled66, 67, 68Nominal

This Week’s Completed MODIS Activities:

Saturday, September 9, 2000

None

Sunday, September 10, 2000

None

Monday, September 11, 2000

None

Tuesday, September 12, 2000

2000/256 14:12:44 Real-Time SMIR VDET/ITWK to 79/110
2000/256 14:38:44 ATC Load OA16 - SDSM Double Screened
2000/256 16:17:38 ATC Load OA15 - SDSM Double Open
2000/256 16:39:03 Real-Time SMIR VDET/ITWK to 110/226 (Nominal Operations)
2000/256 16:45:28 Real-Time Memory Dump of Table 4 (PV gains)
2000/256 18:17:46 Real-Time Set new Gain Values for Band 14H (detectors 1-10)
2000/256 18:19:49 Real-Time Set new Gain Value for Band 25 (detector 1)
2000/256 18:40:37 Real-Time Memory Dump of Table 4 (PV gains)

Wednesday, September 13, 2000

None

Thursday, September 14, 2000

None

Friday, September 15, 2000

2000/259 19:42:22 Real-Time Memory Dump, XIO Read Memory Fault Status
2000/259 19:43:30 Real-Time Memory Dump, XIO Read Memory Fault Address
2000/259 19:44:41 Real-Time Memory Dump, XIO Read OAS Register

This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Saturday, September 16, 2000

None

Sunday, September 17, 2000

None

Monday, September 18, 2000

2000/262 05:05:12 ATC Load - Set Formatter to Night Rate 00:03:15 early (for 50/50 SSR buffer)
2000/262 05:35:22 ATC Load - Set Formatter to Day Rate (for Lunar Roll)
2000/262 05:35:24 ATC Load - Set SCIABNORM Flag to ABNORM
2000/262 05:35:26 ATC Load - PC DC Restore OFF
2000/262 05:35:28 ATC Load - PV DC Restore OFF
2000/262 05:35:30 ATC Load - Sector Rotation to -3072 (EA to SV and OBCs)
2000/262 05:41:44 ATC Load - Sector Rotation to 0 (Normal)
2000/262 05:41:46 ATC Load - PC DC Restore ON
2000/262 05:41:48 ATC Load - PV DC Restore ON
2000/262 05:41:50 ATC Load - Set SCIABNORM Flag to NORM
2000/262 05:41:52 ATC Load - Set Formatter to Night Rate (for Lunar Roll)
2000/262 06:01:09 ATC Load - Set Formatter to Day Rate 00:03:15 late (for 50/50 SSR buffer)

Maneuvers:

The next roll maneuver to view the moon through the space view port is scheduled for September 18th (2000-262, 05:29:56 - 05:16:44 GMT).

MODIS Anomalies:

The first formatter error in 3-1/2 weeks was detected at 256-14:12:42. The software patch that is in place prevented a watchdog reset so operations continued nominally. The formatter temperature does not appear to have been abnormal at the time of the error.

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches.

MODIS Telemetry Trends:

The Radiative Cooler intermediate stage temperature has slowly been increasing since approximately day 240. During this time the LWIR focal plane heater margin has been decreasing. The IOT is continuing to trend these telemetry mnemonics.

Non-MODIS Significant Events:

None

Limited Life Item Status:

SRCA 10W Lamp #1: 173.5 of 500 hours
SRCA 10W Lamp #2: 135.9 of 500 hours
SRCA 10W Lamp #3: 145.1 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 556.1 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1411 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements